two cards are drwan in succession from a deck without replacement.what is the probability that both cards are greater than 2 and less than 8
Answer by rothauserc(4718)

3, 4, 5, 6, 7 so we have 5 cards in each suit for a total of 5 * 4 = 20 cards
:
Probability both cards drawn, without replacement, are > 2 and < 8 = (20/52) * (19/51) = 0.1433 approximately 0.14
: